OpenStudy (anonymous):

x^3=3x^2+16 and i'm looking for x :)

OpenStudy (anonymous):

x^3=3x^2+16 x^3-3x^2-16=0

OpenStudy (anonymous):

Now factor the above cubic polynomial..

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

not at all, can you do this ?

OpenStudy (anonymous):

x^3=3x^2+16 x^3-3x^2-16=0 x^3-4x^2+x^2-4x+4x-16=0 x^2(x-4)+x(x-4)+4(x-4)=0 (x-4)(x^2+x+4)=0

OpenStudy (anonymous):

Can u continue from here?

OpenStudy (anonymous):

from (x^2+x+4) is only delta?

OpenStudy (anonymous):

ok, but delta <0, so x=4, right?

OpenStudy (anonymous):

Yep..... x=4 is the only real solution.

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

WELCOME

OpenStudy (anonymous):

:)